# QUOTA: Quantifying Objects with Text-to-Image Models for Any Domain

### v1

@@ 6 Conclusion
In this work, we presented QUOTA, a novel optimization framework for domain-agnostic object quantification in generative text-to-image models. Unlike existing methods that require retraining for each new domain, QUOTA enables accurate object quantification across diverse and unseen domains through a dual-loop meta-learning strategy that optimizes a domain-invariant prompt. By integrating learnable quantification and domain tokens within prompt learning, our approach captures stylistic variations and maintains accuracy, even for object classes not seen during training. We introduced a new benchmark specifically for evaluating domain-generalizable object quantification, allowing rigorous assessment of model adaptability and quantification accuracy across domains. Extensive experiments demonstrate that QUOTA outperforms conventional models in both object quantification accuracy and semantic consistency, setting a new standard for efficient, scalable, and flexible text-to-image generation across any domain.
7 Acknowledgment
This work is financially supported by the Inception Institute of Artificial Intelligence, the University of Amsterdam, the Cisco gift funding, and the allowance Top consortia for Knowledge and Innovation (TKIs) from the Netherlands Ministry of Economic Affairs and Climate Policy.

### v2

@@ 6 Conclusion
We presented QUOTA, a novel optimization framework for domain-agnostic object quantification in text-to-image generation. Unlike existing methods that require retraining for each domain, QUOTA enables accurate quantification across diverse and unseen domains via a dual-loop meta-learning strategy that optimizes a domain-invariant prompt. By introducing learnable quantification and domain tokens into the prompt space, our approach effectively captures stylistic variations and maintains accuracy—even on object classes not seen during training. To facilitate systematic evaluation, we proposed a new benchmark tailored for domain-generalizable object quantification, enabling rigorous assessment of model adaptability and precision. Extensive experiments demonstrated that QUOTA outperforms standard baselines in both count accuracy and semantic alignment, establishing a new standard for flexible and scalable text-to-image generation.
Limitations. Despite its strong cross-domain performance, QUOTA faces challenges in highly stylized or abstract domains where object boundaries are ambiguous. The dual-loop meta-learning introduces additional training overhead due to iterative prompt optimization, though test-time efficiency remains higher than adaptation-based methods such as IoCo [35]. Furthermore, our reliance on external object detectors may introduce counting errors. Prior work [22] has shown that models like CLIP struggle with numerical reasoning, indicating that integrating counting-aware modules could further improve accuracy. Future work will explore stronger object detectors and extend QUOTA to other generative models, enabling applications in fine-grained attribute control and compositional scene generation.
Disclaimer
Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them.
Acknowledgments
This was supported by European Union’s Horizon Europe research and innovation programme under grant agreement number 101214398 (ELLIOT).
Appendix A Performance Across Object Quantities
Figure 6: MAE curves across object quantities for Painting and Cartoon domains. QUOTA significantly reduces error compared to SDXL [23] and IoCo [35], especially in the common range of 5–20 objects.
Table 7 and Table 6 present the MAE results for 19 classes across two target domains: Painting and Cartoon, covering object quantities from 1 to 25. The results provide a detailed comparison between SDXL [23], IoCo [35], and QUOTA.
In the Painting domain, QUOTA achieves the lowest MAE in 20 out of 25 quantities, consistently outperforming both IoCo [35] and SDXL [23]. Similarly, in the Cartoon domain, QUOTA ranks first in 19 out of 25 cases, demonstrating its robustness in counting objects within stylized visual contexts. These findings highlight the strong performance of QUOTA in capturing complex visual variations and maintaining accurate object quantification across diverse classes and quantity ranges, further validating its effectiveness in real-world applications requiring precise object counting.
Figure 6 complements these tables by visualizing the per-count MAE curves for the two target domains. The curves reveal that QUOTA’s largest gains concentrate in the practically most frequent range of 5–20 objects. In this interval, our method reduces the MAE by an average of 42.4% over SDXL [23] and 38.3% over IoCo [35] in the Painting domain, and by 36.0% and 28.8% respectively in the Cartoon domain. Below five objects, all three models already achieve very low errors, offering limited potential for further gains. In contrast, for counts above 20, the MAE remains high for all methods, likely due to increased visual complexity. These results indicate that QUOTA is particularly effective in the count range that matters most for real-world applications, while maintaining comparable performance on simpler cases.
Appendix B Robust Generalization to Unseen Classes
As shown in Figure 7, we train the models using the ’apples’ class in the painting target domain and evaluate their generalization on unseen classes with semantic similarity: ’tomatoes,’ ’oranges,’ and ’strawberries.’ While IoCo [35] performs well on the seen class during training, it fails to generalize effectively to unseen classes, leading to inaccurate quantification across all tested classes. Similarly, SDXL [23] struggles to adapt, producing inconsistent object counts and exhibiting misalignment with textual prompts.
In contrast, our method, QUOTA, exhibits strong generalization, with accurate quantification across unseen classes. This highlights QUOTA ’s ability to adapt to unseen classes and preserve accuracy, even across stylistically distinct domains. These results highlight the effectiveness of QUOTA in addressing the challenges of cross-domain and cross-class generalization.
Figure 7: Generalization to unseen classes. Comparison of object quantification between SDXL [23], IoCo [35], and our QUOTA, evaluated on both trained and unseen classes.